2.16 Consider a hypothetical X-Y ion pair for which the equilibrium interionic spacing and bonding energy values are 0.38 nm2.14 The net potential energy between two adjacent ions, EN, may be represented by the sum of Equations 2.9 and 2.11; that isA 9d oats rsth sSH EA (2.9B ER r (2.11) n2.14 The net potential energy between two adjacent ions, EN, may be represented by the sum of Equations 2.9 and 2.11; that is

2.16 Consider a hypothetical X-Y ion pair for which the equilibrium interionic spacing and bonding energy values are 0.38 nm and -6.13 eV,
2.14 The net potential energy between two adjacent ions, EN, may be represented by the sum of Equations 2.9 and 2.11; that is, B EN= (2.17) r Calculate the bonding energy Eo in terms of the parameters A, B, and n using the following procedure: 1. Differentiate EN with respect to r, and then set the resulting expression equal to zero, because the curve of EN versus r is a minimum at Eo 2. Solve for in terms of A, B, and n, which yields Fo the equilibrium interionic spacing Determine the expression for Eo by substitut ing ro into Equation 2.17.
A 9d oats rsth sSH EA (2.9
B ER r' (2.11) n
2.14 The net potential energy between two adjacent ions, EN, may be represented by the sum of Equations 2.9 and 2.11; that is, B EN= (2.17) r Calculate the bonding energy Eo in terms of the parameters A, B, and n using the following procedure: 1. Differentiate EN with respect to r, and then set the resulting expression equal to zero, because the curve of EN versus r is a minimum at Eo 2. Solve for in terms of A, B, and n, which yields Fo the equilibrium interionic spacing Determine the expression for Eo by substitut ing ro into Equation 2.17.
respectively. If it is known that n in Equation 2.17 has a value of 10, using the results of Problem 2.14, determine explicit expressions for attractive and repulsive energies EA and ER of Equations 2.9 and 2.11
Answer #1

1 -(-tm + EN = (c) dy dy ) A -n- -A G) 2 + B(-) m1B Ju A 2 Mtl-2 A A fer) Eo= EN-fr) Now Eo A EoEo - A () A 13-A Eo= Thies Ho erpresvore xt-y 39mm Eo=-61Sev -6.13 TS-A A Only And eve Vombler B - AXO stove Lo mst proede fo